# Resonances in b-EMRIs: playing the black hole piano

### v1

@@ V Discussion
The resonant response of spinning SMBHs is substantially more difficult to disentangle due to the richer structure of their QNM spectrum. In the Schwarzschild limit, modes with different azimuthal numbers $m$ are degenerate for a given angular index $\ell$, whereas for Kerr BHs this degeneracy is lifted. The resulting denser spectrum makes individual resonances harder to isolate, as illustrated in Fig. 11. On the other hand, QNMs of rapidly spinning SMBHs are less strongly damped, which may lead to sharper resonances if individual modes can be identified. From the perspective of the stellar-mass binary, a spinning SMBH resembles an ellipsoidal piano, with different orientations exciting different notes (c.f. Sec. II). We leave a detailed exploration of these effects for future work, so as not to obscure the main focus of this article, which we believe establishes a solid foundation for the study of resonant QNM excitation by b-EMRIs.
@@
Acknowledgments. We thank Ana Carvalho for producing Figs. 1 and 2. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant no. DNRF184. We acknowledge support by VILLUM Foundation (grant no. VIL37766). VC acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. AL was supported in part by NSF grant AST-2307888 and the NSF CAREER award PHY-2340457. JN was partially funded by Fundação para a Ciência e Tecnologia (FCT), Portugal, through grant no. UID/4459/2025, and by the European Union’s Horizon 2020 research and innovation programme H2020-MSCA-2022-SE through project EinsteinWaves, grant agreement no. 101131233. MvdM acknowledges financial support provided under the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” grant agreement no. GWSky–101167314. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007855 and No 101131233, as well as from Fundação para a Ciência e a Tecnologia under the project 2024.04456.CERN. This work is supported by Simons Foundation International [70] and the Simons Foundation [71] through Simons Foundation grants SFI-MPS-BH-00012593-09 and SFI-MPS-BH-00012593-11.

### v2

@@ V Discussion
The resonant response of spinning SMBHs is substantially more difficult to disentangle due to the richer structure of their QNM spectrum. In the Schwarzschild limit, modes with different azimuthal numbers $m$ are degenerate for a given angular index $\ell$, whereas for Kerr BHs this degeneracy is lifted. The resulting denser spectrum makes individual resonances harder to isolate, as illustrated in Fig. 11. On the other hand, QNMs of rapidly spinning SMBHs are less strongly damped, which may lead to sharper resonances if individual modes can be identified. From the perspective of the stellar-mass binary, a spinning SMBH resembles an ellipsoidal piano, with different orientations exciting different notes (cf. Sec. II). We leave a detailed exploration of these effects for future work, so as not to obscure the main focus of this article, which we believe establishes a solid foundation for the study of resonant QNM excitation by b-EMRIs.
@@
VI Acknowledgments
We thank Ana Carvalho for producing Figs. 1 and 2. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant no. DNRF184. We acknowledge support by VILLUM Foundation (grant no. VIL37766). VC and J.S.S. acknowledge financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. V. C. and J. S. S. also thank the Fundação para a Ciência e Tecnologia (FCT), Portugal, for the financial support to the Center for Astrophysics and Gravitation (CENTRA/IST/ULisboa) through Grant No. UID/PRR/00099/2025 and Grant No. UID/00099/2025. AL was supported in part by NSF grant AST-2307888 and the NSF CAREER award PHY-2340457. JN was partially funded by Fundação para a Ciência e Tecnologia (FCT), Portugal, through grant no. UID/4459/2025, and by the European Union’s Horizon 2020 research and innovation programme H2020-MSCA-2022-SE through project EinsteinWaves, grant agreement no. 101131233. MvdM acknowledges financial support provided under the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” grant agreement no. GWSky–101167314. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007855 and No 101131233, as well as from Fundação para a Ciência e a Tecnologia under the project 2024.04456.CERN. This work is supported by Simons Foundation International [71] and the Simons Foundation [72] through Simons Foundation grants SFI-MPS-BH-00012593-09 and SFI-MPS-BH-00012593-11.